# Rethinking Layer Removal: A Hybrid Pruning Framework Combining Layer Removal and Singular Value Selection for Efficient LLM Compression

### v2

@@ Abstract
Layer removal is an effective technique for compressing large language models (LLMs) by reducing redundancy and improving inference efficiency. However, indiscriminate pruning disrupts representation stability, leading to performance degradation. We propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), which preserves representation-critical singular values to mitigate these effects. Unlike direct layer removal, GRASP leverages gradient-based attribution on a syntax- and semantics-rich dataset to guide the selection of representation-critical singular values. By selectively applying singular value decomposition (SVD) to affected layers, GRASP achieves efficient compression while maintaining representation stability with minimal overhead. Experiments across multiple LLMs show that GRASP consistently outperforms existing compression methods in perplexity and downstream task performance. The source code is available at GRASP.
@@
Rethinking Layer Removal: A Hybrid Pruning Framework Combining Layer Removal and Singular Value Selection for Efficient LLM Compression
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, including language generation, reasoning, and question answering (Brown et al., 2020; Touvron et al., 2023b). However, their massive parameter sizes pose computational and memory challenges, hindering deployment on resource-limited devices (Zhou et al., 2024). To address this, model compression techniques such as quantization (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023), knowledge distillation (Gu et al., 2023; Xu et al., 2024), and pruning(Sun et al., 2024; Ashkboos et al., 2024) have been widely explored. Among these, structured pruning methods, especially layer removal (Men et al., 2024; Yang et al., 2024; Song et al., 2024), have shown promise by leveraging redundancy within layers of LLMs to achieve significant parameter reductions and inference speedup.
Figure 1: Zero-shot performance of GRASP and pure layer removal under 20% compression ratio. Dashed lines represent perplexity on WikiText-2, while bar charts show the percentage of the original model’s average accuracy on seven reasoning datasets.
Despite its simplicity and efficiency, layer removal introduces two main challenges, as shown in Figure 1. First, while task accuracy may remain stable in some cases, the perplexity (PPL) often degrades significantly, reflecting disruptions in the model’s internal representation consistency. This degradation arises from impaired information flow and misaligned intermediate representations (Liang et al., 2024). The sharp increase in PPL suggests that layer removal eliminates critical transformations necessary for maintaining coherent representation structures, potentially limiting the model’s generalization on complex or unseen tasks. Second, the impact of layer removal varies significantly across models. For example, under the same pruning ratio, LLaMA2 (Touvron et al., 2023b) maintains reasonable performance, while LLaMA3 (Dubey et al., 2024) experiences a steep performance drop. This discrepancy suggests that layer removal affects models unevenly, as certain structural transformations are more essential for maintaining stable representations, leading to performance inconsistencies across models.
Our analysis suggests that these issues arise from the removal of transformations that play an essential role in preserving internal consistency and maintaining structured representations. In this paper, we introduce GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel hybrid pruning framework that integrates layer removal with gradient-based singular value selection for efficient LLM compression. Unlike conventional pruning techniques that either remove full layers or apply global SVD compression, GRASP selectively applies singular value decomposition (SVD) only to layers affected by pruning. A key component of GRASP is its use of a small, syntax- and semantics-rich dataset (e.g., WikiText-2), which serves as a representation evaluation dataset to guide the selection of singular values. Instead of relying solely on magnitude-based ranking, GRASP leverages gradient attribution computed from this dataset to dynamically identify and retain singular values critical for preserving language representations. This ensures that despite layer removal, the model retains essential syntactic and semantic structures, minimizing degradation in perplexity and downstream task performance.
We evaluate GRASP on 10 datasets and five models from two distinct LLM families (LLaMA, Mistral). Extensive experiments validate the effectiveness of GRASP. It achieves compression rates comparable to pure layer removal while significantly improving perplexity and task performance across a wide rage of downstream tasks. Notably, GRASP maintains strong robustness across different models without requiring additional training. Furthermore, it achieves strong performance with minimal calibration data—using only 64 samples from a general-purpose, unlabeled dataset, demonstrating the efficiency of our approach. Additionally, post-training compensation further enhances GRASP’s performance, highlighting its adaptability in compressing large language models while preserving their linguistic capabilities.
Figure 2: (A): Overview of GRASP. The figure illustrates how GRASP uses gradient attribution to identify representation-critical singular values and applies SVD to compress redundant layers while preserving representation-relevant transformations. (B): SVD-based LLM compression and (C): Layer removal compression methods
2 Related Work
2.1 Layer Removal-based LLM Compression
Layer removal is a structured pruning technique that eliminates redundant layers within Large Language Models (LLMs), aiming to reduce the model’s computational and memory requirements. Methods such as ShortGPT (Men et al., 2024) introduce a metric called Block Influence to assess the significance of individual layers, enabling efficient one-shot removal of less important layers. SLEB (Song et al., 2024) improves this by employing an iterative pruning strategy, evaluating the importance of each layer based on the current state of the layer-removed LLMs. LaCo (Yang et al., 2024), on the other hand, proposes a gradual compression approach, progressively merging redundant layers from deeper to shallower parts of the network.
Although effective in reducing model size, layer removal disrupts representation coherence, leading to performance degradation and increased perplexity, as analyzed by Liang et al. (2024). To mitigate this, post-training compensation methods have been proposed. Kim et al., 2024 introduced Shortened LLaMA, which employs LoRA (Hu et al., 2022) to restore pruned models’ capabilities. However, LoRA modifies the singular value spectrum, potentially weakening pre-trained features (Shuttleworth et al., 2024). LLM-Streamline (Chen et al., 2024) addresses this by training a lightweight module, such as an FFN or transformer layer, to approximate the removed layers. While effective, these methods impose high computational and data costs, limiting feasibility in resource-constrained settings.
2.2 SVD-based LLM Compression
Singular Value Decomposition (SVD) is widely used in model compression to decompose weight matrices into low-rank structures, typically selecting top-k singular values based on Frobenius norm reconstruction loss. Recent methods have enhanced SVD to reduce LLM compression error. FWSVD (Hsu et al., 2022) incorporates Fisher information to reweight the importance of parameters before applying SVD. ASVD (Yuan et al., 2023) uses activation patterns from a representation evaluation dataset to scale weight matrices, reducing compression-induced activation errors. SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss. Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output features of each layer with low-rank matrices.
GRASP differs from these methods in three key aspects. First, rather than minimizing F2 norm reconstruction loss, GRASP optimizes a representation learning attribution loss, computed on a small representation evaluation dataset to ensure selection aligns with model representation rather than weight approximation. Second, instead of using a top-k singular value selection strategy, GRASP leverages gradient attribution to dynamically retain representation-critical singular values, which are not necessarily the largest in magnitude. Third, while SVD-based pruning is typically applied to all layers, GRASP is selectively used only in pruned layers, ensuring that layer removal does not disrupt critical transformations.
@@ 3 Method
3.1 Overview and Motivation
In this paper, we propose GRASP, a novel hybrid gradient-guided compression framework designed to compress redundant layers in LLMs while preserving representation-critical transformations. Figure 2 provides an overview of GRASP. The method consists of two main steps: (1) identifying redundant layers based on cosine similarity, and (2) selecting and retaining representation-critical singular values and their corresponding singular vectors using gradient-based attribution. Below, we describe each step in detail.
@@ 3.2 Redundant Layer Selection
3.3 Gradient-Guided Singular Value Selection
3.3.1 Representation Learning Attribution Loss (RLAL)
To address this limitation, GRASP introduces gradient-based attribution to evaluate the importance of each singular value based on its contribution to model performance, rather than relying on magnitude alone. Unlike traditional SVD-based methods that select top-$k$ singular values using a reconstruction loss measured by the Frobenius norm ($F_{2}$ norm), GRASP utilizes a Representation Learning Attribution Loss (RLAL), designed specifically for general representation capability modeling.
RLAL is computed on a small, unlabeled representation evaluation dataset derived from corpora rich in syntactic and semantic structures, such as WikiText-2 and similar datasets. These datasets ensure that the retained singular values are those most crucial for maintaining the model’s ability to capture fundamental language structures, including syntax and semantics, even after compression. Unlike task-specific objectives, RLAL focuses on preserving linguistic consistency and general-purpose representation learning rather than optimizing for any particular downstream application.
Mathematically, RLAL is formulated as:
$\mathcal{L}_{\text{RLAL}}=-\sum_{t}\log P(y_{t}\mid x_{\leq t})$ (2)
where $y_{t}$ is the target token at time step $t$, $x_{\leq t}$ denotes the sequence of preceding tokens, and $P(y_{t}\mid x_{\leq t})$ represents the predicted probability of the next token given the context. This loss function aligns with the standard language modeling loss, ensuring that the model learns meaningful representations.
3.3.2 Singular Value Decomposition (SVD)
For a redundant layer’s weight matrix $W_{\text{redundant}}\in\mathbb{R}^{m\times n}$, SVD decomposes $W_{\text{redundant}}$ as:
$W_{\text{redundant}}=U\Sigma V^{T}$ (3)
where $\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{k})$ contains the singular values ($\sigma_{i}\geq 0$), and $U\in\mathbb{R}^{m\times m}$, $V\in\mathbb{R}^{n\times n}$ are orthogonal matrices containing the left and right singular vectors, respectively.
3.3.3 Representation-Critical Singular Value Importance
To assess the representation-critical importance of each singular value, we introduce $\Phi(\sigma_{i})$, the Representation-Critical Singular Value Importance, which quantifies the contribution of $\sigma_{i}$ to the $\mathcal{L}_{\text{RLAL}}$. Using its first-order Taylor expansion, the alteration in the $\mathcal{L}_{\text{RLAL}}$ caused by $\sigma_{i}$ is approximated as:
$\Delta\mathcal{L}_{\text{RLAL}}(\sigma_{i})\approx\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}\Delta\sigma_{i}$ (4)
where $\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}$ represents the gradient sensitivity of $\sigma_{i}$ to the representation learning attribution loss , and $\Delta\sigma_{i}$ is the perturbation applied to $\sigma_{i}$.
The gradient sensitivity is calculated by projecting the gradient of $\mathcal{L}_{\text{RLAL}}$ with respect to $W_{\text{redundant}}$ onto the singular directions:
$\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}=\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$ (5)
where $G=\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial W}$ is the gradient of the $\mathcal{L}_{\text{RLAL}}$ with respect to $W_{\text{redundant}}$, and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ are the left and right singular vectors corresponding to $\sigma_{i}$. The detailed proof can be found in the Appendix A.1.
Integrating the gradient sensitivity with the magnitude of $\sigma_{i}$, we define $\Phi(\sigma_{i})$ as:
$\Phi(\sigma_{i})=\sigma_{i}\cdot\left(\mathbf{u}_{i}^{T}G\mathbf{v}_{i}\right)$ (6)
where $\Phi(\sigma_{i})$ measures the overall contribution of $\sigma_{i}$ to representation performance, considering both its magnitude and alignment with the gradient of the general-purpose model loss. A larger value of $\Phi(\sigma_{i})$ indicates greater importance.
We rank singular values in descending order of representation relevance based on $\Phi(\sigma_{i})$. Only the top-$k$ singular values with the highest $\Phi(\sigma_{i})$ are retained, and the compressed weight matrix is reconstructed as:
$W_{\text{RLAL}}=\sum_{i=1}^{k}\mathbf{u}_{i}\sigma_{i}\mathbf{v}_{i}^{T}$ (7)
This ensures that $W_{\text{RLAL}}$ preserves the transformations most critical to representation stability, achieving efficient compression while maintaining representation consistency.
3.4 Layer-wise Low-Rank Decomposition
GRASP applies a layer-wise backward compression strategy, starting from the last redundant layer and moving sequentially backward. For each redundant layer, weight matrices are compressed using Gradient-Guided Singular Value Decomposition, retaining only the top-$k$ singular values ranked by $\Phi(\sigma_{i})$. This ensures a low-rank approximation that preserves representation-critical transformations.
By processing one layer at a time, this approach significantly reduces memory overhead, avoiding the need to load the entire model into memory. Additionally, the backward order ensures that compressing deeper layers does not disrupt the consistency of earlier layers, preserving the overall stability of the model. The detailed implementation for Layer-wise Low-Rank Decomposition is provided in Algorithm 1 in the Appendix A.2.
@@ 4 Experiments
In this section, we conduct comprehensive experiments to evaluate GRASP from four key perspectives. (1) We first compare our method with several popular pruning-based LLM compression methods to demonstrate its effectiveness (Section 4.2). (2) Next, we compare GRASP with SVD-based LLM compression methods under different compression ratios with different LLMs (Section 4.3). (3) Then, we further evaluate GRASP on the LongBench (Bai et al., 2023) datasets to assess its performance on long-form text processing tasks (Section 4.4). (4) Finally, we analyze the inference speed-up achieved by GRASP (Section 4.5).
@@ 4.1 Experimental Setup
Below we detail the models, benchmarks and implementation details, with more experimental setups provided in Appendix A.3 due to page limit.
Models

### v3

# GRASP: Replace Redundant Layers with Adaptive Singular Parameters for Efficient Model Compression

@@ Abstract
Recent studies have demonstrated that many layers are functionally redundant in large language models (LLMs), enabling model compression by removing these layers to reduce inference cost. While such approaches can improve efficiency, indiscriminate layer pruning often results in significant performance degradation. In this paper, we propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel compression framework that mitigates this issue by preserving sensitivity-aware singular values. Unlike direct layer pruning, GRASP leverages gradient-based attribution on a small calibration dataset to adaptively identify and retain critical singular components. By replacing redundant layers with only a minimal set of parameters, GRASP achieves efficient compression while maintaining strong performance with minimal overhead. Experiments across multiple LLMs show that GRASP consistently outperforms existing compression methods, achieving 90% of the original model’s performance under 20% compression ratio. The source code is available at GRASP.
@@
GRASP: Replace Redundant Layers with Adaptive Singular Parameters for Efficient Model Compression
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, including language generation, reasoning, and question answering (Brown et al., 2020; Touvron et al., 2023b). However, their massive parameter sizes pose computational and memory challenges, hindering deployment on resource-limited devices (Zhou et al., 2024). To address this, model compression techniques such as quantization (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023), knowledge distillation (Gu et al., 2023; Xu et al., 2024), and pruning(Sun et al., 2024; Ashkboos et al., 2024) have been widely explored. Among these, structured pruning methods remove entire components such as neurons or layers to streamline the model, thereby achieving hardware efficiency and inference speedup.
Figure 1: Unlike conventional layer pruning, which either skips redundant layers—often causing moderate performance drops—or replaces them with lightweight modules that require additional training, GRASP (right) retains only the most critical 10% of parameters within the redundant layers, effectively preserving accuracy with minimal overhead.
In this work, we focus on structured layer pruning, which builds on prior findings that certain consecutive layers in large language models (LLMs) are functionally redundant. These findings have inspired approaches that either remove such layers entirely (Men et al., 2024, Yang et al., 2024, Kim et al., 2024) or replace them with lightweight modules (Chen et al., 2024) to reduce inference cost. Although layer removal is simple and computationally efficient, it often results in significant performance degradation. This degradation arises from disrupted information flow and misaligned intermediate representations (Liang et al., 2024), suggesting that layer removal eliminates certain important components that contribute meaningfully to maintaining model performance. Replacing redundant layers with lightweight modules can mitigate this issue to some extent, but such modules are typically randomly initialized, requiring substantial computational resources for training.
In this paper, we propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel compression framework that replaces redundant layers with adaptive singular parameters for efficient LLM compression. Unlike direct layer removal, GRASP exploits the low-rank structure inherent in redundant layers, replacing the redundant layers with only a small subset of parameters while maintaining strong model performance. Specifically, GRASP operates in two key steps: First, it identifies layers suitable for pruning based on the cosine similarity of output hidden states between adjacent layers. Then, instead of relying on magnitude-based heuristics, GRASP leverages gradient attribution derived from a small calibration dataset to adaptively identify and retain the singular values most critical for downstream task performance.
To evaluate the effectiveness of GRASP, We conduct extensive experiments on 19 datasets and 5 models from two distinct LLM families (LLaMA and Mistral). Notably, GRASP achieves strong performance in a training-free setting, requiring no additional optimization. Furthermore, when post-training compensation is applied, only a small number of samples are needed to rapidly restore model performance. This efficiency arises from retaining critical components within redundant layers, rather than relying on randomly initialized replacements. Overall, this paper makes the following contributions:
  • We propose GRASP, a novel training-free compression framework that replaces redundant layers with adaptive singular parameters, leveraging the low-rank structure within LLMs to preserve performance with minimal overhead.
  • We introduce a gradient-based singular value selection mechanism, enabling efficient identification of critical components without relying on magnitude-based heuristics.
  • We conduct extensive experiments across ten datasets and five models from two major LLM families (LLaMA and Mistral), demonstrating that GRASP consistently achieves strong performance under both training-free and low-resource fine-tuning settings.
@@ 2 Method
Figure 1 illustrates the workflow of GRASP. The method consists of two main steps: Identifying redundant layers to be pruned (Sec 2.1) and replacing redundant layers with critical singular components guided by gradient-based attribution (Sec 2.2). Below, we describe each step in detail.
@@ 2.1 Redundant Layer Selection
2.2 Layer Replacement with Adaptive Singular Parameters
Motivation.
GRASP operates under the hypothesis that redundant layers exhibit an inherent low-rank structure, allowing their functionality to be effectively approximated using low-rank matrices. Based on this insight, a straightforward approach is to apply singular value decomposition (SVD) to these layers. However, prioritizing components solely based on singular value magnitude does not necessarily correlate with downstream task performance (Hsu et al., 2022; Hua et al., 2025). To validate this point, we selectively zero out groups of singular values in the weight matrices of large language models and measure their impact on downstream tasks. As shown in Figure 2, we make two key observations: (1) The contribution of a singular value to downstream task performance is not solely determined by its magnitude; smaller values can be crucial for task performance in some cases. (2) Redundant layers exhibit a highly low-rank structure, where only a few singular directions dominate the model performance.
Figure 2: Sensitivity analysis of grouped singular value truncation. While singular values are typically ordered by magnitude, their impact on downstream performance does not follow the same order.
Key Design.
To address this limitation, GRASP introduces gradient-based attribution to evaluate the importance of each singular value based on its contribution to model performance, rather than relying on magnitude alone. Formally, the singular value decomposition of a weight matrix $W\in\mathbb{R}^{m\times n}$ is given by $W=U\Sigma V^{\top}$. Equivalently, $W$ can be represented as the sum of its rank-one components:
$W=\sum_{k=1}^{l}u_{k}\sigma_{k}v_{k}^{\top},$ (2)
where $l=\min(m,n)$, $\sigma_{k}$ denotes the $k$-th singular value, and $u_{k}$, $v_{k}$ are the $k$-th column vectors of the orthogonal matrices $U$ and $V$, respectively. For brevity, we use $\Phi_{k}$ to represent the singular group $\{u_{k},\sigma_{k},v_{k}^{\top}\}$. GRASP estimates the importance of each $\Phi_{k}$ using a small, general-purpose calibration dataset (e.g., WikiText-2), computing a sensitivity-based score that reflects its effect on model performance, given by:
$I(\Phi_{k})=T(\sigma_{k})+\sum_{i=1}^{m}T(u_{k,i})+\sum_{i=1}^{n}T(v_{k,i})$ (3)
where $T(\cdot)$ denotes the estimated change in loss when a certain parameter $\theta$ is zeroed out. This can be approximated by the second-order Taylor series expansion (LeCun et al., 1989), which is defined as:
$T(\theta)=\left|\theta^{\top}\nabla_{\theta}\mathcal{L}+\frac{1}{2}\theta^{\top}H\theta+\mathcal{O}(\|\theta\|^{3})\right|$ (4)
$\mathcal{L}=-\sum_{t}\log P(y_{t}\mid x_{\leq t})$ (5)
where $\nabla_{\theta}\mathcal{L}$ denotes the gradient of the standard language modeling objective function $\mathcal{L}$ with respect to the parameter $\theta$, and $H$ represents the corresponding Hessian matrix. To reduce computational overhead, we omit the second-order term and approximate the importance using only the first-order derivative (Hua et al., 2025, Kim et al., 2024):
$I(\Phi_{k})=\left|\sigma_{k}\frac{\partial\mathcal{L}}{\partial\sigma_{k}}\right|+\sum_{i=1}^{m}\left|u_{k,i}\frac{\partial\mathcal{L}}{\partial u_{k,i}}\right|+\sum_{i=1}^{n}\left|v_{k,i}\frac{\partial\mathcal{L}}{\partial v_{k,i}}\right|$ (6)
The first term in Eq. 6 captures the gradient projection of the loss $\mathcal{L}$ with respect to the singular value $\sigma_{i}$. The detailed proof is provided in Appendix A.1. By aggregating the first-order expansions across all components within each singular group, our method effectively captures the contribution of singular values. A larger value indicates a greater importance of the singular group. This gradient-oriented attribution moves beyond heuristic magnitude-based criteria, enabling performance-aligned importance evaluation. Under the hypothesis that redundant layers exhibit an inherent low-rank structure, we retain only the top-$k$ singular values most critical to model performance based on Eq. 6, and use them to replace the corresponding redundant layer.
2.3 Detailed Implementation of GRASP
GRASP processes the redundant layers in a sequential manner, starting from the final redundant layer and proceeding backward through the network. We summarize the detailed algorithm in Algorithm 1.
Algorithm 1 GRASP: Gradient-based Retention of Adaptive Singular Parameters
1:Model $M$, Calibration set $D$, Retain ratio $r$
2:Compressed model $\tilde{M}$
3:Step1: Redundant Layer Selection:
4:Compute cosine similarity $\text{cos}(H_{i},H_{i+1})$ for all layers
5:Select top-$L$ layers with highest similarity as redundant
6:Step 2: Gradient-Guided Compression:
7:for each redundant layer $l$ (in reverse order) do
[…]
9:         SVD: $W=U\Sigma V^{\top}$
10:         Compute importance $I(\Phi_{k})$ for each $\Phi_{k}=\{u_{k},\sigma_{k},v_{k}^{\top}\}$ via Eq. 6
11:         Keep top-$r\%$ singular groups, reconstruct $\tilde{W}$
12:     end for
13:end for
14:return $\tilde{M}$
@@ 3 Experiments
In this section, we conduct comprehensive experiments to evaluate GRASP from three key perspectives. (1) We first compare our method with existing pruning-based LLM compression approaches to demonstrate its effectiveness (Section 3.2). (2) Next, we analyze the inference speed-up achieved by GRASP (Section 3.3). (3) Finally, we investigate the factors influencing our approach by performing ablation studies on the choice of calibration datasets and pruning strategies. (Section 3.4)
@@ 3.1 Experimental Setup
Below we detail the models, benchmarks, baselines and implementation details used in our experiments, with more experimental setups provided in Appendix A.2.
Models.